OpenStudy (anonymous):

On Monday, Lucy’s bank balance was –$164. On Tuesday, her bank balance was –$17. What was the change in Lucy’s bank balance from Monday to Tuesday? $

OpenStudy (anonymous):

if u help i will try to give metal

OpenStudy (texaschic101):

-164 - (-17) = -164 + 17 = - 147 she started with a negative balance of 164...she then added 147 to make her new balance be -17.

OpenStudy (anonymous):

sorry all ready got it, it was $147

OpenStudy (texaschic101):

lol...sorry that I was too late

OpenStudy (anonymous):

it is all cool man

OpenStudy (anonymous):

here you can help with 2 more

OpenStudy (anonymous):

is that cool

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

it will not let me post it i will just put up a new one k

OpenStudy (texaschic101):

sure thing